How do I find the total perimeter of 2 squares, one with an area of the square root of 5 and the other with a square root of 125?
Answer by stanbon(57967)

How do I find the total perimeter of 2 squares, one with an area of the square root of 5 and the other with a square root of 125?
----
Area of a square = (side)^2
--------
1st square:
area = sqrt(5)
Solve s^2 = sqrt(5)
s = 5^(1/4)
Perimeter = 4*5^(1/4)
-----------------------------
2nd sqruare:
area = sqrt(125) = 5sqrt(5)
Solve s^2 = 5sqrt(5)
s = 5^(2/4)*5^(1/4) = 5^(3/4)
Perimeter = 4*5^(3/4)
===========================
Cheers,
stan H.